OpenStudy (hari5719):

2 ^y=8 i mean 2 power y equal to 8 whats is y

OpenStudy (hari5719):

@abb0t

OpenStudy (hari5719):

@tayromo

OpenStudy (hari5719):

@ganeshie8

OpenStudy (hari5719):

@dan815

OpenStudy (aum):

8 = 2 x 2 x 2 2^y = 8 = 2^3 2^y = 2^3 y = ?

OpenStudy (dan815):

2*2*2=2^3=8

OpenStudy (hari5719):

wat is y

OpenStudy (dan815):

lol

OpenStudy (dan815):

are you trolling hari!

OpenStudy (aum):

\(\large 2^y = 2^3\) What is y?

OpenStudy (hari5719):

3

OpenStudy (aum):

Yes!

OpenStudy (hari5719):

i was trolling

OpenStudy (dan815):

2 to the some number = 8 2^3 was 8 so 3 is that some number

OpenStudy (dan815):

ok